OpenStudy (winsteria):

MEDAL AND FAN A bag contains 5 red cubes and 7 white cubes. Celinda will select one cube, put it in her pocket, and then select a second cube. Which expression gives the probability that she will select two red cubes? A. 5/12 . 5/12 B. 5/12 . 4/12 C. 5/12 . 5/11 D. 5/12 . 4/11

OpenStudy (anonymous):

D

OpenStudy (cwrw238):

what is the probability that she picks a red on the first draw?

OpenStudy (anonymous):

Im right

OpenStudy (cwrw238):

yes you are right but plz dont just give the answer

OpenStudy (cwrw238):

as there are 5 reds and 7 white - a total of 12 cubes then P( pick e red) = 5/12

OpenStudy (cwrw238):

after one red is put into the pocket we have 11 cubes left of which 4 are red so P( red) = 4/11

OpenStudy (cwrw238):

so we have 2 events - the 2 draws from the bag - and they are independent events . Independent because one event cannot have any effect on the other. In this case the individual probabilities are multiplied so the required probability = 5/12 * 4/11

OpenStudy (winsteria):

Thank You

OpenStudy (cwrw238):

yw